OpenStudy (anonymous):

How do I solve: 2a-3b+c=10 2a-2b-2c=2 a+3b+2c=-1

OpenStudy (agentnao):

Multiply the entire first equation by -2. Now subtract the second equation from the first. (The c variable will be eliminated) Isolate "a" in the third equation and plug into the difference from the step above.

OpenStudy (anonymous):

thanks

OpenStudy (agentnao):

I hope that helps. If you still don't get it, let me know and I'll show you step by step.

OpenStudy (anonymous):

c is not eliminated when multiplying the entire equation by -2.

OpenStudy (anonymous):

multiply by -2 and then subtract the ENTIRE second equasion from the ENTIRE first equasion, do you know how to set this up?

OpenStudy (anonymous):

kind of...

OpenStudy (anonymous):

this will make "c" cancel out.

OpenStudy (anonymous):

right 6a-8b=24

OpenStudy (alexwee123):

use matrices ! and reduce to row echelon form

OpenStudy (anonymous):

@alexwee123 wait, what?

OpenStudy (alexwee123):

its only if you know matrices then it's so much simpler

OpenStudy (anonymous):

+8b to both sides and divide by 6, a = 4 + 4/3b plug in A to 6a-8b=24 to solve for b plug in a and b into any of the original equasions to solve for x

OpenStudy (anonymous):

Im only in Alg 2. They only taught us 2x2 matrices.

OpenStudy (alexwee123):

well what did you do to solve out the 2x2?

OpenStudy (anonymous):

the b's cancel out when I plug A into 6a-8b=24..

OpenStudy (anonymous):

alright then. plug in the "a" we found into a+3b+2c=-1, and then get either variable isolated.

OpenStudy (anonymous):

regardless as to what variable you choose, (say you choose to isolate b) you then use this to plug into any of the three equasions, knowing what "a" and "b" equal. You now have only one variable, "c", which you solve for. Knowing "c", you plug in this number for the equasion 4 + 4/3b+3b+2c=-1 and solve for "b". Knowing "b", you plug in this number in any equasion, along with the plugging in of "c". You now can find "a"!!!

OpenStudy (alexwee123):

read this if you have time http://www.math.fsu.edu/~bellenot/class/f08/lalab/other/rref2.pdf